OpenStudy (anonymous):

a line passes through the points R(-1,-5) and S(-3,-1). write the equation of the line through the point A(4,0)that is parallel to RS.

OpenStudy (anonymous):

a.y=-2x-2 b.y=-2x+8 c.y=2x-8 d. y=2x

OpenStudy (anonymous):

@Darknight200 @humphriescody @keke_luvs_u @Owlfred @midhun.madhu1987 @walter20 @taylor1216

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

c

OpenStudy (anonymous):

It's B

OpenStudy (anonymous):

is it b or c? @humphriescody @taylor1216 @keke_luvs_u

OpenStudy (anonymous):



OpenStudy (anonymous):

I said C.

OpenStudy (anonymous):

c